OpenStudy (anonymous):

if a+b+c=2, what does -8c-8a-8b equal???

OpenStudy (anonymous):

take -8 common an put value of a+b+c

OpenStudy (anonymous):

hint \[-8(a+b+c)=?\]

OpenStudy (anonymous):

oh ok -16

OpenStudy (anonymous):

how did you know to do that????? I have been trying to rearrange a+b+c=2 to find the value of the variables and then plug those in

OpenStudy (anonymous):

\[-8c-8a-8b \] by the distributive property this is \[-8(c+a+b)=-8(a+b+c)\]

OpenStudy (anonymous):

since \[a+b+c=2\] you know \[-8(a+b+c)=-8\times 2=-16\]

OpenStudy (anonymous):

brilliant